# ValuesRAG: Enhancing Cultural Alignment Through Retrieval-Augmented Contextual Learning

### v2

@@ 4.3 Retrieval-based Values Alignment
Role-Assignment 0.5598 0.7493 0.4770 0.6048 0.4747 0.5262 0.5653
Few-Shot Learning 0.5225 0.6376 0.4559 0.5782 0.4194 0.4758 0.5149
@@
Model Num(K) EVS GSS CGSS ISD LAPOP Africa Avg.
@@
Table 3: Accuracy scores across six regional datasets for the ablation study (Section 5.3.2). Num($K$) ($k\in\{1,3,5,10\}$) indicates the number of demographic summaries retrieved. Bold text indicates the best performance, underlined text indicates the second-best performance.
5 Experiments
5.1 Setup
Models Used.
We utilize GPT-4o-mini Achiam et al. (2023) and Gemini-1.5-Flash Team et al. (2024) for our generation tasks, which are accessed via APIs. We set the temperature parameter of these models to $0.7$ to achieve a balance between coherence and creativity. For the retrieval task, we employed the E5 (base) model Wang et al. (2022) [^4] https://huggingface.co/intfloat/e5-base, which generates embeddings and retrieves the top 100 most relevant summaries of values based on cosine similarity. Additionally, we utilize the GTE-multilingual-reranker-base model Zhang et al. (2024) [^5] https://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base for reranking. This reranking model refines the selection process and ensures that the top-$k$ most relevant summaries are selected from the initial top-100 retrievals.
Baseline Methods and Implementation.
Our baseline methods include: (1) Zero-shot inference, (2) the role-assignment-only approach Tao et al. (2024), (3) a few-shot learning method Choenni and Shutova (2024), and (4) a hybrid method that combines both (1) and (2).
Specifically, for the role-assignment baseline, we use the same demographic summaries as in ValuesRAG to ensure fairness by assigning roles based on demographic information from the survey data. For the few-shot method, we follow the approach outlined in the previous work, where we randomly select five examples from the test set as prompts. The hybrid method combines both strategies, assigning roles based on demographic summaries and augmenting the prompts with five randomly selected few-shot examples from the test set. [^6] Detailed prompts are provided in Appendix B.
Additionally, We use ValuesRAG with $k=3$ retrieved summaries—chosen to provide a good balance between retrieval diversity and contextual relevance. [^7] Detailed analysis of varying $k$ and its implications on model performance is provided in Section 5.3.1.
Evaluation.
We utilize accuracy as the primary evaluation metric by converting multiple-choice responses into a binary format for consistency and simplicity. This is achieved by categorizing responses into two distinct groups based on contrasting answer patterns, effectively capturing agreement and disagreement. This approach aligns with the structure of values-related questions, such as those using a Likert Scale.
5.2 Experimental Analysis
In our experiments, we compare our method, ValuesRAG, with four baseline methods: zero-shot inference, role-assignment, few-shot learning, and a hybrid approach that combines role-assignment and few-shot learning, as shown in Table 2.
We find that the role-assignment method generally surpasses both zero-shot and few-shot approaches. By grounding the agent’s responses in a clearly defined demographic context, it ensures more consistent performance. Yet, role assignment can sometimes lead to overly narrow representations when demograhpic roles are interpreted stereotypically. Meanwhile, few-shot learning can incorporate example-driven context, but its limited number of prompts may not consistently address the intricate ways individuals’ belifs diverge within similar social settings. As a result, it struggles to generalize to the multifaceted nature of human values, particularly when faced with unexpected or complex cultural scenarios. The hybrid method, which merges role assignment and few-shot prompts, does offer a partial improvement in contextual diversity. However, it remains insufficient for capturing the full spectrum of nuances that can arise from overlapping demographic factors and idiosyncratic personal persepectives.
@@ 5.2 Experimental Analysis
5.3 Ablation Study
We conduct two ablation studies to analyze the configuration and robustness of ValuesRAG., in Section 5.3.1, we vary the number of retrieved summaries ($k$) to examine how retrieval depth affects the model’s performance. Second, in Section 5.3.2, we isolate the effect of using only values-based generation.
5.3.1 5.3.1 Varying the Number of Retrieved Summaries
To further investigate the role of retrieval depth in ValuesRAG, we vary the number of retrieved summaries, $k\in\{1,3,5,10\}$. Table 3 presents the performance of ValuesRAG for three models across multiple datasets. We observe that while $k=1$ may not provide enough diversity, increasing $k$ beyond 3 can sometimes introduce less relevant information.
Moreover, $k=3$ consistently achieves significantly better accuracy compared to higher values of $k$, highlighting its effectiveness and achieving the best balance between contextual relevance and model latency; these ensure both effectiveness and efficiency. This configuration not only enhances retrieval accuracy but also maintains low computational overhead, underscoring the strength of ValuesRAG in real-world applications where speed and performance are equally crucial. Hence, we choose $k=3$ as the default retrieval number for conducting our baseline experiments, ensuring optimal performance across various models and datasets.
5.3.2 5.3.2 Impact of Values-Only Generation
To validate the robustness of ValuesRAG, we perform an ablation study using only values context augmented generation, thereby excluding the impact of demographic summaries on the model’s performance. We use the WVS validation set—separated from the training data, which served as the retrieval corpus (as outlined in Section 3)—to evaluate the models. Table 4 presents a comparison of our method, using only summaries of values, against four baseline methods. Notably, ValuesRAG consistently outperforms all baselines across this validation data, achieving the highest accuracy despite relying exclusively on the values summaries.
@@ 5.3 Ablation Study
Table 4: Accuracy comparison between baseline methods and our proposed Values Augmented Generation method using the WVS validation set. Bold text indicates the best performance, underlined text indicates the second-best performance.
@@
This result confirms the effectiveness and robustness of the values-augmented generation approach. ValuesRAG leverages structured values summaries to generate contextually rich and culturally aligned responses. Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined QA examples or demographic anchors. The results demonstrate the framework’s scalability and adaptability, effectively mitigating biases and generating culturally coherent outputs with minimal dependence on external context.
6 Conclusion
We present ValuesRAG, an innovative Retrieval-Augmented Generation framework that revolutionizes cultural values alignment through its comprehensive approach to values context analysis. While existing approaches rely on pre-trained knowledge with fixed demographic labels, often leading to oversimplified generalizations and biases, or few-shot prompting techniques that struggle with generalization due to limited correlations across value dimensions, ValuesRAG overcomes these weaknesses by dynamically retrieving and integrating granular demographic data with detailed individual-level cultural profiles. Our extensive evaluations demonstrate that ValuesRAG achieves an accuracy improvement of up to 21% over state-of-the-art baselines. By combining adaptive retrieval mechanisms with in-context learning and reranking strategies, ValuesRAG captures complex cultural nuances, reduces biases, and ensures contextually aligned responses. This structured design enables ValuesRAG to bridge the gap between generic LLM capabilities and the demands of culturally sensitive applications, providing a scalable and robust solution for real-world use cases.
7 Limitation
@@
Although our baseline comparisons indicate that ValuesRAG generally delivers superior performance compared to alternative methods, it does not always guarantee an exact match to individual’s true values. Since we rely on the WVS dataset to summarize individual profiles, there can be mismatches when these summaries are applied to other test sets. In future work, we plan to explore more adaptive retrieval strategies that can better align with novel datasets, as well as investigate how integrating additional fine-tuning with retrieval-augmented generation may further refine each agent’s contextual accuracy.
Appendix A Appendix A. Dataset Description
A.1 WVS
As shown in Figure 3, the WVS dataset conducts surveys in over 120 countries worldwide, covering most regions and countries across the globe, making it suitable for constructing a documents corpus for Retrieval Augmented Generation (RAG). It is also representative in terms of demographic characteristics. We select the most recent WVS survey (Wave 7) to ensure the dataset is as close as possible to the contemporary world.
Figure 3: Distribution of Country in WVS dataset. The WVS dataset is the most extensive values survey dataset in terms of coverage.
According to the classification in the codebook released by WVS, the values-related questions it contains can be divided into 13 topics, as shown in Table 5, covering most dimensions of values, allowing for a comprehensive and accurate measure of each respondent’s values.
Topic Count
Social Values, Norms, Stereotypes 45
Happiness and Wellbeing 11
Social Capital, Trust and Organizational Membership 47
Economic Values 6
Perceptions of Corruption 9
Perceptions of Migration 10
Perceptions of Security 21
Index of Postmaterialism 6
Perceptions about Science and Technology 6
Religious Values 12
Ethical Values 23
Political Interest and Political Participation 35
Political Culture and Political Regimes 25
A.2 Test Datasets
To validate the robustness of our method, we selected datasets from six different regions as test sets. These test sets include both values-related questions and demographic characteristics. The demographic characteristics are used to generate summaries, serving as retrieval targets for RAG. The values-related questions are utilized as test questions to calculate accuracy (ACC).
EVS
GSS
CGSS
ISD
To ensure that our experiment covers as much of the world’s population as possible, we made efforts to include India within the scope of our test set. However, we were unable to obtain data from several government surveys in India, so we used data published by the Pew Research Center instead. The Pew Research Center’s India Survey Dataset [^11] https://www.pewresearch.org/dataset/india-survey-dataset is a comprehensive resource that captures the perspectives of 29,999 Indian adults on various aspects of society, including religious beliefs and practices, identity, nationalism, and societal tolerance. Conducted through face-to-face interviews between November 17, 2019, and March 23, 2020, the survey encompassed participants from diverse religious backgrounds, such as Hindus, Muslims, Sikhs, Christians, Buddhists, Jains, and others. This dataset covers 33 values-related questions and 23 demographic characteristics.
LAPOP
Africa
Appendix B Appendix B. Prompts Used
This section provides the prompts designed for various components of the ValuesRAG, including prompts for performing question answering tasks, as well as for generating values and demographic summaries.
Appendix C Appendix C. Case Study
Example 1 Question Is premarital sex always wrong, almost always wrong, sometimes wrong, or not wrong at all? Original Demographics Summary The respondent is a 63-year-old male born in the United States, of White descent, who speaks English at home. He completed 12th grade and holds a high school diploma, while his spouse has the same educational background. They are currently married and live in a household of two, with no children. The respondent identifies as middle class and reported a family income of \$90,000 to \$109,999 last year, with his own earnings falling within the same range. He practices Protestantism and is affiliated with the United Presbyterian Church in the U.S.A. His mother has completed two years of college, while his father finished high school. The respondent’s demographic background reflects a stable family life with no financial challenges reported. Retrieved Summaries: Retrieved Text 1
Demographics
The respondent is a 60-year-old White, non-Hispanic male born in the United States, where he speaks English at home. He was raised in a family where both parents were also born in the U.S., and he holds a bachelor’s degree, signaling a level of educational attainment typical of the upper middle class. He is currently employed full-time in a professional and technical role within the private business sector and is the chief wage earner in his household, which consists of one person. The respondent is divorced and has two children. He identifies as a Protestant and has reported saving money over the past year, rating his household income as relatively high on a scale of one to ten.
Values
This person’s values center on family, personal empowerment, and social equity, emphasizing a balanced life encompassing work, political engagement, and community well-being. They prioritize financial responsibility, health, and individual liberties while advocating equality and rejecting traditional gender roles. They approach trust cautiously, emphasize accountability in governance, and hold nuanced views on immigration and moral integrity. They believe in science and technology’s benefits and maintain personal faith amid diverse beliefs. Politically active, they support democratic principles and civil rights while voicing dissatisfaction with current institutions. Overall, they blend progressive values focused on social justice, environmental sustainability, and the power of citizen engagement in politics and public life.
Retrieved Text 2
Demographics
The respondent is a 63-year-old White, non-Hispanic male born in the United States, where he speaks English at home. He is a citizen of the U.S., with both parents also being born in the country. He is married and lives in a household of two people, including himself and his spouse, who is retired. He has two children and both he and his spouse completed 12th grade, earning a high school diploma. He works full-time in a semi-skilled occupation within the private sector and identifies as belonging to the middle class, reporting that his household income places them in the third income tier. The respondent aligns with the Protestant religion and describes their financial situation as stable.
Values
This individual values family and work as life’s center, while moderately engaging in politics. They emphasize independence, hard work, tolerance, and determination in children, embrace progressive stances on gender and social issues, and place little importance on religious conformity. Their trust is selective, relying heavily on family, showing moderate faith in others, and opposing corruption, while valuing economic equality and environmental concerns. They uphold science and technology, favor personal freedom over security, and maintain moral integrity and accountability. Strongly engaged in civic matters, they are committed to democracy, civil rights, and active political participation, yet remain cautious about government roles. Overall, their values balance familial dedication, personal agency, ethical living, and a progressive yet pragmatic outlook on social and political issues.
Retrieved Text 3
Demographics
The respondent is a 65-year-old white male, born in the United States, …
Example 2 Question How many children to have is a personal matter, and the government should not interfere. Do you agree? Original Demographics Summary The respondent is a 27-year-old male from China of Han ethnicity, who speaks Chinese at home. He holds a postgraduate degree and has never engaged in any work for economic income. Currently, he lives alone without any additional household members and identifies as belonging to the middle class, describing his household’s economic status as average in his local area. He does not practice any religion and reports no personal income in the past year. Retrieved Summaries: Retrieved Text 1
Demographics
The respondent is a 28-year-old male of Chinese ethnicity, born in the country where he currently resides. He speaks Standard Chinese (Mandarin/Putonghua/Guoyu) at home and identifies as part of the working class. He has attained a Bachelor’s degree and is employed full-time in a clerical role within the private sector. He lives in a household of four, including his own parents, and is currently married but has no children. Both he and his spouse have a similar educational background and are employed full-time. His family’s financial situation indicates that they have spent some savings and borrowed money over the past year. He does not belong to any religious denomination and has reported his household income as being in the third income level out of ten.
Values
This individual holds a complex set of values that place a significant emphasis on family, work, and personal responsibility, while also promoting income equality and valuing hard work as a means of success. They have a moderate stance on political engagement, expressing skepticism toward certain institutions while maintaining high trust in organizations like the police and government. Their views reflect a critical perspective on issues such as immigration and gender roles, where they value cultural diversity but prioritize native citizens for job opportunities. They exhibit a commitment to environmental protection, scientific advancements, and ethical living, while acknowledging their non-religious identity. Overall, their values convey a blend of personal autonomy, accountability, and a pragmatic view on governance, demonstrating a preference for democratic principles intertwined with the belief in the importance of social order and national security.
Retrieved Text 2
Demographics
The respondent is a 29-year-old male of Chinese descent, and speaks other Chinese dialects at home. He has attained a bachelor’s degree and works full-time in the sales sector within a private business or industry. Living alone, he is single and has no children. His family background includes parents with upper secondary education. Socially, he identifies with the working class and describes his household’s financial status as just getting by, placing his household income in the fourth income group on a scale of ten. He does not identify with any religious denomination.
Values
This person exhibits a set of values that emphasize the importance of family, hard work, and personal responsibility, reflecting a traditional view of societal roles and a commitment to instilling these qualities in children. They hold a pragmatic and somewhat secular worldview, expressing low interest in politics and religion while maintaining a positive attitude towards health, personal agency, and economic stability. They advocate for income equality and environmental sustainability, demonstrating concern for social equity and opposing corruption, while also valuing trust primarily within familial ties. Additionally, their perspective on immigrants is cautiously appreciative, and they prioritize freedom and economic progress, alongside a recognition of the complexities of modern ethical and moral issues. Overall, this individual’s values highlight a balance between familial loyalty, economic and environmental responsibilities, personal freedom, and a discerning approach to broader societal structures.
Retrieved Text 3
Demographics
The respondent is a 25-year-old male of Chinese ethnicity, who speaks Standard Chinese languages, …

### v3

@@ 2. Temporal proximity:
Role-Assignment  (2024) 0.5598 0.7493 0.4770 0.6048 0.4747 0.5262 0.5653
Few-Shot Learning  (2024) 0.5225 0.6376 0.4559 0.5782 0.4194 0.4758 0.5149
@@
Model Num(K) EVS GSS CGSS ISD LAPOP Africa Avg. Accuracy
@@
Table 4: Accuracy scores across six regional datasets for the ablation study: Varying the Number of Retrieved Summaries. Num($K$) ($k\in\{1,3,5,10\}$) indicates the number of demographic summaries retrieved. Bold text indicates the best performance, underlined text indicates the second-best performance.
Experimental Analysis
In our experiments, we compare ValuesRAG with four baseline methods: (1) zero-shot inference, (2) role-assignment, (3) few-shot learning, and (4) a hybrid approach that combines role-assignment and few-shot learning. Following previous work, as detailed in the Evaluation Metrics section, we specifically evaluate model performance using accuracy. All responses are binarized based on contrasting answer patterns to ensure consistency across different values-related questions. Experimental results are shown in Table 3.
We find that the role-assignment method generally surpasses both zero-shot and few-shot approaches. By grounding the agent’s responses in a clearly defined demographic context, it ensures more consistent performance. Yet, role assignment can sometimes lead to overly narrow representations when demograhpic roles are interpreted stereotypically. Meanwhile, few-shot learning can incorporate example-driven context, but its limited number of prompts may not consistently address the intricate ways individuals’ belifs diverge within similar social settings. As a result, it struggles to generalize to the multifaceted nature of human values, particularly when faced with unexpected or complex cultural scenarios. The hybrid method, which merges role assignment and few-shot prompts, does offer a partial improvement in contextual diversity. However, it still remains insufficient for capturing the full spectrum of nuances that can arise from overlapping demographic factors and idiosyncratic personal persepectives.
@@
Ablation Study
We conduct two ablation studies to analyze the configuration and robustness of ValuesRAG. First, we vary the number of retrieved summaries ($k$) to examine how retrieval depth affects the model’s performance. We subsequently isolate the effect of using only values-based generation.
Varying the Number of Retrieved Summaries
To quantify how the number of retrieved summaries $k$ impacts ValuesRAG’s performance, we evaluate $k\in\{1,3,5,10\}$. Table 4 reports accuracy on six regional datasets for both GPT-4o-mini and Gemini 1.5 Flash.
For GPT-4o-mini, increasing $k$ from 1 to 3 boosts accuracy on five of six datasets, rising from an average of 0.6078 to 0.6195, and yields the best scores on GSS (0.7781), CGSS (0.5387), and Africa (0.5953). Although $k=5$ further improves EVS (0.6052) and ISD (0.7016), it degrades GSS and CGSS, dropping the average to 0.6174. At $k=10$, performance falls across most datasets (average 0.6074), indicating that an excessive number of summaries dilute relevance. Similarly, for Gemini 1.5 Flash, $k=3$ attains the highest overall accuracy (0.5977), with top scores on EVS (0.5869), CGSS (0.5337), ISD (0.6789), and LAPOP (0.4705). While $k=5$ slightly edges out $k=3$ on GSS (0.7690) and Africa (0.5498), it lowers EVS and ISD and yields a lower average (0.5959). Retrieval at $k=10$ further declines.
These results reveal a clear trade-off: $k=1$ constrains diversity, whereas $k>3$ introduces marginally relevant or noisy summaries and increases latency. Consequently, $k=3$ achieves the best balance between contextual breadth, accuracy gains, and computational cost. We thus adopt $k=3$ as our default retrieval depth across all experiments.
Impact of Values-Only Generation
To validate the robustness of ValuesRAG, we perform an ablation study using only values context augmented generation, thereby excluding the impact of demographic summaries on the model’s performance. We use the WVS validation set: separated from the training data, which served as the retrieval corpus (as outlined in Section: Datasets), to evaluate the models. Table 5 presents a comparison of our method, using only summaries of values, against four baseline methods. Notably, ValuesRAG consistently outperforms all baselines across this validation data, achieving the highest accuracy despite relying exclusively on the values summaries.
@@
Table 5: Accuracy comparison between baseline methods and Values Augmented Generation method using the WVS validation set. Bold text indicates the best performance, underlined text indicates the second-best performance.
Compared to zero‐shot inference (0.6176 for GPT-4o-mini; 0.6041 for Gemini), Values-Only achieves 0.6894 with GPT-4o-mini and 0.6583 with Gemini, demonstrating structured value context provides substantially richer guidance than unconstrained prompts. Against the role-assignment (0.6747 for GPT-4o-mini; 0.6505 for Gemini), Values-Only attains higher scores as well, confirming that our generated value summaries capture more nuanced cultural patterns than static demographic labels. Similarly, Values-Only outperforms few-shot learning (0.6359 for GPT-4o-mini; 0.6086 for Gemini) and the hybrid method (0.6670 for GPT-4o-mini; 0.6354 for Gemini), further illustrating even without explicit examples or demographic anchors, our values-driven prompts produce more context-aligned predictions.
@@
These results confirm the effectiveness and robustness of the values-augmented generation approach. ValuesRAG leverages structured values summaries to generate contextually rich and culturally aligned responses. Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined QA examples or demographic anchors. The results demonstrate the framework’s scalability and adaptability, effectively mitigating biases and generating culturally coherent outputs with minimal dependence on external context.
Case Study
To further illustrate the effectiveness and practical utility of ValuesRAG, we present two qualitative case studies derived from the GSS (United States) and CGSS (China) datasets. These case studies clearly demonstrate how our method dynamically forms a culturally and demographically relevant retrieval corpus, enhances values alignment, and consistently outperforms baseline approaches by providing richer, context-sensitive reasoning for the generation process.
Figure 4: Case Study (United States): Cultural Attitudes Toward Premarital Sex. The original demographic profile (a 63-year-old Protestant male from the U.S.) is complemented by multiple retrieved summaries, enabling the LLM to reason with contextual sensitivity, avoiding stereotypes and enhancing values alignment.
Case 1: Cultural Attitudes Toward Premarital Sex (United States)
The first case (Figure 4) involves a 63-year-old middle-class Protestant male from the U.S., responding to the question: “Is premarital sex always wrong, almost always wrong, sometimes wrong, or not wrong at all?” ValuesRAG constructs the retrieval corpus by embedding demographic information and retrieving topically relevant profiles with similar backgrounds (e.g., age, ethnicity, education, and religious affiliation). Specifically, the top three retrieved profiles closely align demographically with the original individual yet introduce nuanced ideological variations. Each retrieved document not only reflects the target individual’s demographic features but also articulates distinctive value orientations, highlighting the contextual breadth and depth of the retrieved corpus.
This carefully constructed retrieval corpus significantly enhances values alignment by providing multiple representative and nuanced perspectives, allowing the LLM to consider various contextually valid stances rather than reinforcing stereotypes typically found in static demographic-based methods. Compared to role-assignment or few-shot approaches, which struggle to represent this nuanced cultural diversity, ValuesRAG dynamically integrates subtle but critical distinctions across retrieved summaries. As a result, the generated response reflects realistic and culturally plausible reasoning that respects both traditional values and evolving societal norms, showcasing the superiority of ValuesRAG in handling culturally sensitive and complex value judgments.
Figure 5: Case Study (China): Family Planning Preferences. The retrieved summaries facilitate nuanced reasoning around personal autonomy and governmental roles, effectively capturing contemporary social shifts in attitudes toward family planning.
Case 2: Family Planning Preferences (China)
The second case (Figure 5) examines a 27-year-old secular male with postgraduate education from urban China addressing the question: “How many children to have is a personal matter, and the government should not interfere. Do you agree?” In this scenario, ValuesRAG retrieves profiles of individuals with closely matching demographics—such as age, education, urban residence, etc. Each retrieved summary reflects values related to personal agency, social responsibility, and family planning, with slight variations in economic status or professional roles. The constructed retrieval corpus is well-aligned with the original profile and relevant to the thematic context of the prompt.
By incorporating these diverse yet demographically coherent value profiles, ValuesRAG enables the LLM to generate responses that account for a variety of socially grounded perspectives on family decision-making. In contrast to static prompting methods that may overlook such intra-cultural nuances, ValuesRAG captures shifts in value expression across population segments in a context-aware manner. This case shows how the method’s dynamic retrieval mechanism enhances cultural alignment by offering broader coverage of relevant value orientations within the target demographic.
Conclusion
We propose ValuesRAG, a novel framework designed to advance cultural values alignment through context-aware reasoning. In contrast to prior methods that rely on fixed demographic labels or limited few-shot prompting, ValuesRAG dynamically retrieves and integrates contextually rich value summaries using adaptive retrieval, reranking, and In-Context Learning (ICL). Our contributions extend beyond the framework design to include a comprehensive evaluation across diverse, culturally and geographically representative test datasets. Extensive experiments demonstrate that ValuesRAG consistently outperforms existing approaches, effectively capturing complex cultural nuances, reducing biases, and generating contextually aligned responses. Through its structured and adaptive design, ValuesRAG bridges the gap between general-purpose LLMs and the demands of culturally sensitive applications, offering a scalable and robust solution for real-world deployment.
Limitation
@@ Conclusion
Although our baseline comparisons indicate that ValuesRAG generally delivers superior performance compared to alternative methods, it does not always guarantee an exact match to an individual’s true values. Specifically, since our method relies on dynamically retrieved summaries from a predefined corpus (WVS), mismatches may occur when summaries derived from one dataset are applied to different or novel datasets. Future research should therefore explore more adaptive retrieval strategies capable of precisely aligning with novel datasets and investigate how integrating additional fine-tuning steps within retrieval-augmented generation frameworks could further enhance each agent’s contextual accuracy and generalizability across diverse value domains.
Ethical Considerations
While ValuesRAG effectively mitigates the common stereotype reinforcement seen in static methods, its reliance on demographic features inherently carries potential ethical risks related to profiling, bias, and fairness. If deployed in sensitive or high-stakes contexts such as public policy or persuasive communication, improper handling of demographic information could unintentionally perpetuate stereotypes or biases, adversely affecting vulnerable groups. We emphasize that our framework is intended not to reinforce specific demographic-driven stereotypes or biases, but rather to surface contextually relevant values explicitly for scrutiny.
Researchers and practitioners utilizing ValuesRAG must proactively examine and address ethical implications surrounding demographic-based retrieval. It is critical to integrate ValuesRAG into broader evaluative frameworks designed for continuous monitoring of potential societal impacts, ensuring its responsible deployment. Additionally, future work should explicitly investigate the ethical dimensions of demographic profiling within retrieval-augmented methods, further promoting fairness, accountability, and transparency in AI-driven cultural alignment tasks.